solve by lagrange's method z(p-q) =z^2+(x+y)^2


1
Expert's answer
2021-09-13T05:43:19-0400

Standard pde is Pp+Qq=R,

where p="\frac{ \partial z}{\partial x}" and q="\frac{ \partial z}{\partial y}" .

Comparing the given pde with the standard pde.


We get, P=z, Q=-z, R=z2+(x+y)2


By lagrange method,

"\frac{dx}{P}=\frac{dy}{Q}=\frac{dz}{R}\\
\frac{dx}{z}=\frac{dy}{-z}=\frac{dz}{z^2+(x+y)^2}\\"

By taking first equation,

"\frac{dx}{z}=\frac{dy}{-z}\\
dx=-dy\\
x=-y+c_1\\
x+y=c_1"

By taking second equation,

"\frac{dy}{z}=\frac{dz}{z^2+(x+y)^2}\\
dy=\frac{z}{z^2+c_1^2}dz 
[ \space \text{from first equation}]\\
dy=\frac{1}{2}\frac{2z}{z^2+c_1^2}dz"

Integrating both side, we get

"y=\frac{1}{2}ln(z^2+c_1^2)+c_2\\
y-ln\sqrt{(z^2+(x+y)^2)}=c_2"

Therefore, the solution is given by

"F(x+y, y-ln\sqrt{(z^2+(x+y)^2)})=0"